Ask
your own question, for FREE!
Mathematics
9 Online
The medical records for a class of 28 children show whether they had previously had measles or mumps. The records show 22 have had measles, 13 have had measles and mumps, and 27 have had measles or mumps. If one child from the class is selected at random, determine the probability that he or she has had: part a: measles part b: mumps but not measles part c: neither mumps nor measles.
Still Need Help?
Join the QuestionCove community and study together with friends!
Can help with a and c a. 22/28 of class had measles, so this is probability 78.6% c. 1/28 had neither so this is probability 1/28 = 3.6% b. having mumps but not measles? 27 had mumps or measles, 22 had measles, leaving 5 who had mumps only? 5/28 = 17.9%. Adds up to 28/28, so perhaps this is right.
Can't find your answer?
Make a FREE account and ask your own questions, OR help others and earn volunteer hours!
Join our real-time social learning platform and learn together with your friends!
Join our real-time social learning platform and learn together with your friends!
Latest Questions
VERNEISH:
Two objects are moving along separate linear paths where each path is described by position, d, and time, t.
Maci:
what are the best stock apps for people who want to start investing?
Twaylor:
genuienly how do i get better self control i keep watcing youtube and theres this
IzzyCookie:
|dw:1792074056138:dw| Uhm, I don't know where to ask this, but I'm in college welding at my community college, and I need help learning vertical welding for
thehearken:
not sure if this is considered spam, im sorry if it is, but does anyone have tips
4 hours ago
1 Reply
0 Medals
4 days ago
3 Replies
2 Medals
1 week ago
3 Replies
0 Medals
2 days ago
9 Replies
1 Medal
1 week ago
2 Replies
1 Medal